Metallic magnesium can be made by the electrolysis of molten MgCl2.

(a) What mass of Mg is formed by passing a current of 8.70 A through molten MgCl2 for 3.60 days? The unbalanced chemical reaction representing this electrolysis is shown below.

MgCl2 Mg + Cl2

__________ g of Mg is formed by this electrolysis.

(b) How many minutes are needed to plate out 20.00 g of Mg from molten MgCl2 using 8.78 A current?

__________- minutes are needed.
Answer #1

Answer:

(A) 340.83 g

(B) 301 minutes

explanation:

Answen-Pant (a) We fhove beten given, the cuYrent whlth basses through the Solutton 8-70 A While, th time ot cuYYert tlow 3.6continue..

we put the values in the given formula,

1.64 mole = charge /96485 c

Charge = 1.64*96485 = 158790.37

Again, since we have to calculate the time ,we use the simple formula

Q = i * t, Q= charge, i= current , t = time

Or t = q/i = 158790.37/8.78

t = 18085.46 second,

converting it to minutes by deviding by 60.

t = 18085.46 /60 minutes = 301.42 minutes

please upvote